# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ 4.2.1 Intuitionistic non-deterministic valuations
We define the function $h:For(\Omega)\rightarrow For(\Sigma)$ as follows:
@@
    $h(\neg\alpha)=\neg f_{\Box}(\alpha)$;
  3. iii)
[…]
We call $h$ as the semi-translation function.
Lemma 4.5.
For every $\alpha\in For(\Omega)$, $f_{\Box}(\alpha)=\Box(h(\alpha))$.
Proof.
Let $\alpha$ be an intuitionistic formula. We proceed by case analysis.
Case $\alpha=p_{n}$
In this case, $f_{\Box}(\alpha)=f_{\Box}(p_{n})=\Box p_{m}$. Also, $\Box(h(\alpha))=\Box(h(p_{n}))=\Box p_{m}$. Note that $n=m$ in both cases. Hence, $f_{\Box}(p_{n})=\Box(h(p_{n}))$.
Case $\alpha=\neg\beta$
By Definition 4.4, $\Box(h(\neg\beta))=\Box(\neg f_{\Box}(\beta))$. Moreover, by Definition 4.2, $f_{\Box}(\neg\beta)=\Box(\neg f_{\Box}(\beta))$. Hence, $f_{\Box}(\neg\beta)=\Box(h(\neg\beta))$.
[…]
By Definition 4.4, we have that for each ##\##, $\Box(h(\beta\#\gamma))=\Box(f_{\Box}(\beta)\#f_{\Box}(\gamma))$. Moreover, by Definition 4.2, $f_{\Box}(\beta\#\gamma)=\Box(f_{\Box}(\beta)\#f_{\Box}(\gamma))$. Therefore, $f_{\Box}(\beta\#\gamma)=\Box(h(\beta\#\gamma))$.
Remark 4.6.
By Lemma 4.5 and Remark 4.3, $v(f_{\Box}(\alpha))=\Box^{\mathcal{A}_{S4}}(v(h(\alpha)))$, for every $\alpha\in For(\Omega)$.
Definition 4.7.
(Truth function) Let $v$ be a valuation function in $\mathcal{M}_{S4}$ and $\alpha\in For(\Omega)$. We define the function $\tilde{v}:For(\Omega)\rightarrow V\times V$ as follows:
$\tilde{v}(\alpha)\triangleq\langle v(h(\alpha));v(f_{\Box}(\alpha))\rangle$
We will say that $\tilde{v}$ is induced by $v$ and will call $\tilde{v}$ as the truth function.
By Remark 4.6, we get $\tilde{v}(\alpha)=\langle v(h(\alpha));\Box^{\mathcal{A}_{S4}}(v(h(\alpha)))\rangle$. This motivates the following:
Definition 4.8 (Intuitionistic truth values).
The set of intuitionistic truth values is given by
$V_{IPL}=\left\{\langle c;\Box^{\mathcal{A}_{S4}}(c)\rangle\mid c\in\mathcal{V}\right\}=\{\mathbf{F},\mathbf{U},\mathbf{T}\}$
where $\mathbf{F}=\langle 0;0\rangle$, $\mathbf{U}=\langle 1;0\rangle$ and $\mathbf{T}=\langle 2;2\rangle$.
Therefore, the truth function 4.7 is a function $\tilde{v}:For(\Omega)\rightarrow\{\mathbf{F},\mathbf{U},\mathbf{T}\}$.
[…]
Definition 4.9 (Intuitionistic multioperations).
Let $\langle a,b\rangle,\langle c,d\rangle\in V_{IPL}$. Define the intuitionistic multioperations as follows:
  1. 1.
    $\neg^{\prime\mathcal{A}_{IPL}}(\langle a;b\rangle)=\left\{\langle c;\Box^{\mathcal{A}_{S4}}(c)\rangle\mid c\in\neg^{\mathcal{A}_{S4}}(b)\right\}$
  2. 2.
    #$\#^{\prime\mathcal{A}_{IPL}}(\langle a;b\rangle,\langle c;d\rangle)=\left\{\langle e;\Box^{\mathcal{A}_{S4}}(e)\rangle\mid e\in\#^{\mathcal{A}^{\prime}_{S4}}(b,d)\right\}$
[…]
Definition 4.10.
The intuitionistic Nmatrix over signature $\Omega$ is given by $\mathcal{M}_{IPL}=\langle V_{IPL},\mathcal{D}_{IPL},\mathcal{O}_{IPL}\rangle$, where $\mathcal{D}_{IPL}=\{\mathbf{T}\}$, $\mathcal{O}_{IPL}=\{\neg^{\mathcal{A}_{IPL}},\rightarrow^{\mathcal{A}_{IPL}},\lor^{\mathcal{A}_{IPL}},\land^{\mathcal{A}_{IPL}}\}$ and each $o\in\mathcal{O}_{IPL}$ is defined in Table 3.
$\neg^{\mathcal{A}_{IPL}}$
@@
$\rightarrow^{\mathcal{A}_{IPL}}$ F U T
[…]
$\lor^{\mathcal{A}_{IPL}}$ F U T
[…]
$\land^{\mathcal{A}_{IPL}}$ F U T
[…]
Table 3: Tables for #$\#^{\mathcal{A}_{IPL}}$.
Lemma 4.11.
Let #$\#\in\Omega$. Then, #$\#^{\mathcal{A}_{IPL}}=\#^{\prime\mathcal{A}_{IPL}}$. That is, the multioperators in Definition 4.9 coincide with the ones in Definition 4.10.
@@
It is straightforward to verify that, for every $z,w\in V_{IPL}$:
  1. 1.
    $\neg^{\mathcal{A}_{IPL}}(z)=\neg^{\prime\mathcal{A}_{IPL}}(z)$; and
  2. 2.
[…]
Definition 4.12.
[…]
Definition 4.13.
Definition 4.14.
Let $Val(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$, and let $Val_{+}(\mathcal{M}_{IPL})$ be the set of valuations $v$ in $Val(\mathcal{M}_{IPL})$ such that $v(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
Definition 4.15.
(Level valuations for IPL) Let $Val(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$. We define the level $\mathcal{L}^{IPL}_{k}$, where $k\in\mathbb{N}$, as follows:
  1. i)
    $\mathcal{L}_{0}^{IPL}=Val_{+}(\mathcal{M}_{IPL})$;
  2. ii)
    $\mathcal{L}^{IPL}_{k+1}=\{v\in\mathcal{L}^{IPL}_{k}\mid\forall\alpha\in For(\Omega)_{k+1}$, if $v(\alpha)=\mathbf{U}$ then there exists $w\in\mathcal{L}^{IPL}_{k}$ such that $w(\alpha)=\mathbf{F}$ and: $v(\beta)=\mathbf{T}$ implies $w(\beta)=\mathbf{T}$ for every $\beta\in For(\Omega)\}$.
Definition 4.16.
The set of level valuations in $\mathcal{M}_{IPL}$ is defined as the intersection of every level valuation:
$\mathcal{L}_{IPL}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{IPL}_{k}$
Definition 4.17 (RNmatrix for IPL).
From now on, $H_{IPL}$ will denote the standard Hilbert calculus for IPL over $\Omega$, given by the following axioms, and with Modus Ponens (MP) as the only inference rule:
$\begin{array}[]{ll}\textsf{(Ax1)}&\alpha\to(\beta\to\alpha)\\[5.69054pt] \textsf{(Ax2)}&(\alpha\to(\beta\to\gamma))\to((\alpha\to\beta)\to(\alpha\to\gamma))\\[5.69054pt] \textsf{(Ax3)}&\alpha\to(\beta\to(\alpha\land\beta))\\[5.69054pt] \textsf{(Ax4)}&(\alpha\land\beta)\to\alpha\\[5.69054pt] \textsf{(Ax5)}&(\alpha\land\beta)\to\beta\\[5.69054pt] \textsf{(Ax6)}&\alpha\to(\alpha\lor\beta)\\[5.69054pt] \textsf{(Ax7)}&\beta\to(\alpha\lor\beta)\\[5.69054pt] \textsf{(Ax8)}&(\alpha\to\gamma)\to((\beta\to\gamma)\to((\alpha\lor\beta)\to\gamma))\\[5.69054pt] \textsf{(Ax9)}&(\beta\to\alpha)\to((\beta\to\neg\alpha)\to\neg\beta)\\[5.69054pt] \textsf{(Ax10)}&\alpha\to(\neg\alpha\to\beta)\\[5.69054pt] \end{array}$
The consequence relation of $H_{IPL}$ will be denoted by $\vdash_{IPL}$. Observe that derivations in $H_{IPL}$ from premises are defined as usual in Hilbert calculi. That is, premises can be used in the derivations in the same way as axioms.
The proof of soundness of $H_{IPL}$ w.r.t. the RNmatrix $\mathcal{R}(\mathcal{M}_{IPL})$ will be postponed to Subsection 4.2.2. We will concentrate now on the completeness proof.
@@
Recall from Subsection 3.2 the notion of $\varphi$-saturated set in L. If $\Delta$ is a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$ (that is, IPL) then: $\alpha\in\Delta$ implies that $\neg\alpha\notin\Delta$; $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; $\alpha\to\beta\in\Delta$ and $\alpha\in\Delta$ implies that $\beta\in\Delta$; and $\beta\in\Delta$ implies that $\alpha\to\beta\in\Delta$.
@@
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$ (that is, IPL). The valuation associated to $\Delta$ in IPL is the function $v_{\Delta}:For(\Omega)\to V_{IPL}$ defined inductively as follows (here, $p$ is a propositional variable):

### v4

@@ 8.2 Level valuations for $\mathcal{M}_{IPL}$
Let $Val^{nd}(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$, and let $Val^{nd}_{+}(\mathcal{M}_{IPL})$ be the set of valuations $v$ in $Val^{nd}(\mathcal{M}_{IPL})$ such that $v(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
Definition 8.6.
(Level valuations for $\mathbf{IPL}$) The set of $n$th-level valuations $\mathcal{L}^{IPL}_{n}$ for $\mathbf{IPL}$, where $n\in\omega$, is defined as follows:
@@
    $\mathcal{L}^{IPL}_{n+1}=\{v\in\mathcal{L}^{IPL}_{n}\mid\forall\alpha\in For(\Omega)$, if $v(\alpha)=\mathbf{U}$ then there exists $w\in\mathcal{L}^{IPL}_{n}$ such that $w(\alpha)=\mathbf{F}$ and, for every $\beta\in For(\Omega)$: $v(\beta)=\mathbf{T}$ implies $w(\beta)=\mathbf{T}\}$.
The set of level valuations in $\mathcal{M}_{IPL}$ is defined as the intersection of the sets of $n$th-level valuations, for $n\geq 0$:
$\mathcal{L}_{IPL}=\bigcap^{\infty}_{n\geq 0}\mathcal{L}^{IPL}_{n}$
Definition 8.7 (RNmatrix for IPL).
8.3 Soundness and Completeness of $\mathbf{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$
From now on, LI will stand for a standard sequent calculus for IPL over $\Omega$. In fact, the sequent calculus LI presented here is obtained from Gentzen’s system LJ (see, for instance, [30, Chapter 1]) with minor modifications. Concretely, we consider sequents formed by pairs of finite multisets of formulas instead of pairs of sequences (which allows to eliminate the exchange rule), and by adding contexts to the axiom (which allows to eliminate the left weakening rule). Certain rules can also be simplified in this setting. Clearly, both calculi are equivalent.
By a sequent we mean an expression $\Gamma\Rightarrow\Delta$ such that $\Gamma$ and $\Delta$ are (non simultaneously empty) finite multisets, and $\Delta$ contains at most one formula. The calculus LI is given by the following axiom and inference rules:
Axiom
$(Ax)\,\,\alpha,\Gamma\Rightarrow\alpha$
Structural Rules
$(contr\Rightarrow)\,\displaystyle\frac{\alpha,\alpha,\Gamma\Rightarrow\Delta}{\alpha,\Gamma\Rightarrow\Delta}\hskip 56.9055pt(\Rightarrow w)\,\displaystyle\frac{\Gamma\Rightarrow}{\Gamma\Rightarrow\alpha}$
$(cut)\,\displaystyle\frac{\Gamma\Rightarrow\alpha\hskip 14.22636pt\alpha,\Gamma\Rightarrow\Delta}{\Gamma\Rightarrow\Delta}$
Logical Rules
@@
$\mbox{($\neg\Rightarrow$)}\,\,\displaystyle\frac{\Gamma\Rightarrow\alpha}{\neg\alpha,\Gamma\Rightarrow}\hskip 42.67912pt\mbox{($\Rightarrow\neg$)}\,\,\displaystyle\frac{\alpha,\Gamma\Rightarrow}{\Gamma\Rightarrow\neg\alpha}$
The consequence relation of LI will be denoted by $\vdash_{IPL}$. Hence, for any set $\Gamma\cup\{\alpha\}\subseteq For(\Omega)$, $\Gamma\vdash_{IPL}\alpha$ iff there exists a finite set $\Gamma_{0}\subseteq\Gamma$ such that the sequent $\Gamma_{0}\Rightarrow\alpha$ is derivable in LI. The logic generated by LI, with consequence relation $\vdash_{IPL}$, is precisely $\mathbf{IPL}$.
Remark 8.8.
Being a presentation of $\mathbf{IPL}$, the deduction metatheorem (DM) holds for the consequence relation $\vdash_{IPL}$ associated to LI:
$(DM)\ \ \Gamma,\alpha\vdash_{IPL}\beta\ \ \mbox{ iff }\ \ \Gamma\vdash_{IPL}\alpha\to\beta$
for every set of formulas $\Gamma\cup\{\alpha,\beta\}$. The validity of (DM) can be checked directly by using ($Ax$) and the rules for implication. Indeed, the ‘only if’ part follows by ($\Rightarrow\to$). For the ‘if’ part, it can be observed that $\Gamma_{0},\alpha,\alpha\to\beta\Rightarrow\beta$ is provable in LI for every finite multiset $\Gamma_{0}$ and formulas $\alpha$ and $\beta$: this is a consequence of applying ($\to\Rightarrow$) to the axioms $\Gamma_{0},\alpha\Rightarrow\alpha$ and $\Gamma_{0},\alpha,\beta\Rightarrow\beta$. From this, the result follows from the hypothesis $\Gamma_{0}\Rightarrow\alpha\to\beta$ (for a finite set $\Gamma_{0}\subseteq\Gamma$) and (cutcut).
Definition 8.9.
Let $\Pi$ be a derivation of a sequent $\Gamma\Rightarrow\Delta$ in LI (that is, $\Pi$ is a tree of sequents with root $\Gamma\Rightarrow\Delta$ and axioms as leaves). The depth of $\Pi$ is its height (as a tree). If a sequent $\Gamma\Rightarrow\Delta$ admits a derivation in LI of depth $k$ (for $k\geq 0$), we write $\Gamma\vdash_{IPL}^{k}\Delta$.
The proof of soundness of LI w.r.t. the RNmatrix $\mathcal{R}(\mathcal{M}_{IPL})$ requires to establish, in a precise way, soundness of derivations of depth $k$ w.r.t. valuations of level $k$.
Definition 8.10.
Let $k\geq 0$ and $v\in\mathcal{L}_{k}^{IPL}$. We say that $v$ satisfies a formula $\alpha$, written as $v\models\alpha$, if $v(\alpha)=\mathbf{T}$. If $\Gamma$ is a non-empty (multi)set of formulas, we say that $v$ satisfies $\Gamma$, denoted by $v\models\Gamma$, if $v\models\gamma$ for every $\gamma\in\Gamma$. We say that $v$ satisfies a sequent $\Gamma\Rightarrow\Delta$, denoted by $v\models\Gamma\Rightarrow\Delta$, if the following holds:
(1) $\Delta=\emptyset$. In this case, $v\not\models\Gamma$, that is, $v(\beta)\in\{\mathbf{U},\mathbf{F}\}$ for some $\beta\in\Gamma$; or
(2) $\Delta=\{\gamma\}$, for some formula $\gamma$. In this case, $v\models\gamma$ whenever $v\models\Gamma$.
A sequent $\Gamma\Rightarrow\Delta$ is said to be valid in $\mathcal{L}_{k}^{IPL}$, denoted by $\Gamma\models_{k}\Delta$, if it is satisfied by any $v\in\mathcal{L}_{k}^{IPL}$.
Proposition 8.11.
For every $k\geq 0$ and every sequent $\Gamma\Rightarrow\Delta$: if $\Gamma\vdash_{IPL}^{k}\Delta$, then $\Gamma\models_{k}\Delta$.
@@
By induction on $k$.
Base $k=0$: Suppose that $\Gamma\vdash_{IPL}^{0}\Delta$. Then $\Gamma\Rightarrow\Delta$ is an instance of the axiom ($Ax$), that is: $\Gamma\Rightarrow\Delta=\alpha,\Gamma^{\prime}\Rightarrow\alpha$ for some multiset $\Gamma^{\prime}$. Let $v\in\mathcal{L}_{0}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\alpha\}$. In particular, $v\models\alpha$, hence $v\models\Gamma\Rightarrow\Delta$. From this, $\Gamma\models_{0}\Delta$.
Inductive step: Suppose that, for any sequent $\Gamma\Rightarrow\Delta$, if $\Gamma\vdash_{IPL}^{j}\Delta$ then $\Gamma\models_{j}\Delta$, for every $j\leq k$, for a given $k\geq 0$ (Induction Hypothesis, IH). Let $\Gamma\Rightarrow\Delta$ be a sequent such that $\Gamma\vdash_{IPL}^{k+1}\Delta$. If $\Gamma\Rightarrow\Delta$ is an instance of the axiom, it can be proven as above that $\Gamma\models_{k+1}\Delta$. Suppose now that $\Gamma\Rightarrow\Delta$ was obtained from other sequents by applying a rule of LI in a derivation of depth $k+1$. We have the following cases:
($contr\Rightarrow$): Suppose that $\Gamma\Rightarrow\Delta=\alpha,\Gamma^{\prime}\Rightarrow\Delta$ is obtained from $\alpha,\alpha,\Gamma^{\prime}\Rightarrow\Delta$ by applying rule ($contr\Rightarrow$). Then $\alpha,\alpha,\Gamma^{\prime}\vdash_{IPL}^{k}\Delta$ and so $\alpha,\alpha,\Gamma^{\prime}\models_{k}\Delta$, by (IH). But the latter clearly implies that $\alpha,\Gamma^{\prime}\models_{k}\Delta$ and so $\alpha,\Gamma^{\prime}\models_{k+1}\Delta$, given that $\mathcal{L}_{k+1}^{IPL}\subseteq\mathcal{L}_{k}^{IPL}$.
($\Rightarrow w$): Suppose that $\Gamma\Rightarrow\Delta=\Gamma\Rightarrow\alpha$ is obtained from $\Gamma\Rightarrow$ by ($\Rightarrow w$). Then $\Gamma\vdash_{IPL}^{k}$ and so, by (IH), $\Gamma\models_{k}$. This means that, for every $v\in\mathcal{L}_{k}^{IPL}$, $v\not\models\Gamma$ and so $v\models\Gamma\Rightarrow\alpha$. From this it follows that $\Gamma\models_{k+1}\Delta$, given that $\mathcal{L}_{k+1}^{IPL}\subseteq\mathcal{L}_{k}^{IPL}$.
(cutcut): Assume that $\Gamma\Rightarrow\Delta$ follows from $\Gamma\Rightarrow\alpha$ and $\alpha,\Gamma\Rightarrow\Delta$ by (cutcut) rule. Then $\Gamma\vdash_{IPL}^{k}\alpha$ and $\alpha,\Gamma\vdash_{IPL}^{k}\Delta$ and so, by (IH), $\Gamma\models_{k}\alpha$ and $\alpha,\Gamma\models_{k}\Delta$. Suppose first that $\Delta=\emptyset$. Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma$; in particular, $v\in\mathcal{L}_{k}^{IPL}$ and so $v\models\Gamma\Rightarrow\alpha$. From this it follows that $v\models\alpha$. But this is impossible, given that $v\models\Gamma\cup\{\alpha\}\Rightarrow$, by (IH). From this, $v\not\models\Gamma$ for every $v\in\mathcal{L}_{k+1}^{IPL}$. That is, $v\models\Gamma\Rightarrow$, for every $v\in\mathcal{L}_{k+1}^{IPL}$. Now, suppose that $\Delta=\{\gamma\}$ for some formula $\gamma$, and let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma$. Once again, since $v\models\Gamma\Rightarrow\alpha$ we infer that $v\models\alpha$. But then $v\models\Gamma\cup\{\alpha\}$ and so $v\models\gamma$, given that $v\models\Gamma\cup\{\alpha\}\Rightarrow\gamma$. This means that $v\models\Gamma\Rightarrow\Delta$, for every $v\in\mathcal{L}_{k+1}^{IPL}$. That is, $\Gamma\models_{k+1}\Delta$.
($\vee\Rightarrow$): Assume that $\Gamma\Rightarrow\Delta=\alpha\vee\beta,\Gamma^{\prime}\Rightarrow\Delta$ follows from $\alpha,\Gamma^{\prime}\Rightarrow\Delta$ and $\beta,\Gamma^{\prime}\Rightarrow\Delta$ by ($\vee\Rightarrow$). By (IH), $\alpha,\Gamma^{\prime}\models_{k}\Delta$ and $\beta,\Gamma^{\prime}\models_{k}\Delta$. Suppose first that $\Delta=\emptyset$. By hypothesis, if $v\in\mathcal{L}_{k+1}^{IPL}$ is such that $v\models\Gamma^{\prime}$, then $v\not\models\alpha$ and $v\not\models\beta$, and so $v\not\models\alpha\vee\beta$. This means that $v\models\Gamma^{\prime}\cup\{\alpha\vee\beta\}\Rightarrow$, for every $v\in\mathcal{L}_{k+1}^{IPL}$. Now, if $\Delta=\{\gamma\}$ for some formula $\gamma$, let $v\in\mathcal{L}_{k+1}^{IPL}$. By hypothesis, if $v\models\Gamma^{\prime}$ and $v\models\alpha$, then $v\models\gamma$; and if $v\models\Gamma^{\prime}$ and $v\models\beta$, then $v\models\gamma$. From this, if $v\models\Gamma^{\prime}$ and $v\models\alpha\vee\beta$ then, by definition of $\mathcal{M}_{IPL}$, either $v\models\alpha$ or $v\models\beta$ and so, by hypothesis, $v\models\gamma$. That is, $v\models\Gamma\Rightarrow\Delta$ for every $v\in\mathcal{L}_{k+1}^{IPL}$ and so $\Gamma\models_{k+1}\Delta$.
($\Rightarrow\vee_{i}$): It is enough to observe that, if either $v\models\alpha$ or $v\models\beta$ then $v\models\alpha\vee\beta$, by definition of $\mathcal{M}_{IPL}$.
($\wedge\Rightarrow$): Suppose that $\Gamma\Rightarrow\Delta=\alpha\land\beta,\Gamma^{\prime}\Rightarrow\Delta$ follows from $\alpha,\beta,\Gamma^{\prime}\Rightarrow\Delta$ by ($\wedge\Rightarrow$). Assume first that $\Delta=\emptyset$. Then, if $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}$ it follows that $v\not\models\{\alpha,\beta\}$, since $\mathcal{L}_{k+1}^{IPL}\subseteq\mathcal{L}_{k}^{IPL}$ and $\alpha,\beta,\Gamma^{\prime}\models_{k}\Delta$ by (IH). From this, either $v\not\models\alpha$ or $v\not\models\beta$. By definition of $\mathcal{M}_{IPL}$, $v\not\models\alpha\land\beta$. That is, $v\not\models\Gamma^{\prime}\cup\{\alpha\land\beta\}$ for every $v\in\mathcal{L}_{k+1}^{IPL}$, i.e., $\alpha\wedge\beta,\Gamma^{\prime}\models_{k+1}$. In turn, if $\Delta=\{\gamma\}$ for some formula $\gamma$, let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\alpha\land\beta\}$. By definition of $\mathcal{M}_{IPL}$, $v\models\Gamma^{\prime}\cup\{\alpha,\beta\}$ and so $v\models\gamma$, by (IH) and the fact that $v\in\mathcal{L}_{k}^{IPL}$. That is, $v\models\Gamma^{\prime}\cup\{\alpha\land\beta\}\Rightarrow\Delta$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
($\Rightarrow\wedge$): It is enough to observe that, if both $v\models\alpha$ and $v\models\beta$ then $v\models\alpha\wedge\beta$, by definition of $\mathcal{M}_{IPL}$.
($\to\Rightarrow$) Suppose that $\Gamma\Rightarrow\Delta=\alpha\to\beta,\Gamma^{\prime}\Rightarrow\Delta$ is obtained from $\Gamma^{\prime}\Rightarrow\alpha$ and $\beta,\Gamma^{\prime}\Rightarrow\Delta$ by ($\to\Rightarrow$). By (IH), $\Gamma^{\prime}\models_{k}\alpha$ and $\beta,\Gamma^{\prime}\models_{k}\Delta$. Assume first that $\Delta=\emptyset$. Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\alpha\to\beta\}$. Then, $v\in\mathcal{L}_{k}^{IPL}$ such that $v\models\Gamma^{\prime}$ and so $v\models\alpha$, by hypothesis. Since $v\models\alpha\to\beta$ then $v\models\beta$, by definition of $\mathcal{M}_{IPL}$. That is, $v\in\mathcal{L}_{k}^{IPL}$ is such that $v\models\Gamma^{\prime}\cup\{\beta\}$, which contradicts the fact that $\beta,\Gamma^{\prime}\models_{k}$. From this, $v\not\models\Gamma^{\prime}\cup\{\alpha\to\beta\}$ for every $v\in\mathcal{L}_{k+1}^{IPL}$, which means that $\alpha\to\beta,\Gamma^{\prime}\models_{k+1}$. Now, suppose that $\Delta=\{\gamma\}$ for some formula $\gamma$. Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\alpha\to\beta\}$. By reasoning as above, $v\in\mathcal{L}_{k}^{IPL}$ is such that $v\models\Gamma^{\prime}\cup\{\beta\}$. But then $v\models\gamma$, by hypothesis. This implies that $v\models\Gamma^{\prime}\cup\{\alpha\to\beta\}\Rightarrow\gamma$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
($\Rightarrow\to$): Assume that $\Gamma\Rightarrow\Delta=\Gamma\Rightarrow\alpha\to\beta$ is obtained from $\alpha,\Gamma\Rightarrow\beta$ by ($\Rightarrow\to$). By (IH), $\alpha,\Gamma\models_{k}\beta$. Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma$. Suppose that $v(\alpha\to\beta)=\mathbf{F}$. By definition of $\mathcal{M}_{IPL}$, $v(\alpha)=\mathbf{T}$ and $v(\beta)\in\{\mathbf{U},\mathbf{F}\}$. That is, $v\models\Gamma\cup\{\alpha\}$ but $v\not\models\beta$, which contradicts the fact that $\alpha,\Gamma\models_{k}\beta$, given that $v\in\mathcal{L}_{k}^{IPL}$. Hence, $v(\alpha\to\beta)\neq\mathbf{F}$. Suppose now that $v(\alpha\to\beta)=\mathbf{U}$. By Definition 8.6 of level valuations, there exists $w\in\mathcal{L}_{k}^{IPL}$ such that $w(\alpha\to\beta)=\mathbf{F}$ and $w\models\Gamma$. By reasoning as above, $w(\alpha)=\mathbf{T}$ and $w(\beta)\in\{\mathbf{U},\mathbf{F}\}$. But then $w\models\Gamma\cup\{\alpha\}$ and $w\not\models\beta$, which contradicts the fact that $\alpha,\Gamma\models_{k}\beta$. From this, it follows that $v(\alpha\to\beta)=\mathbf{T}$. That is, $v\models\Gamma\Rightarrow\alpha\to\beta$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
($\neg\Rightarrow$): Assume that $\Gamma\Rightarrow\Delta=\neg\alpha,\Gamma^{\prime}\Rightarrow$ is obtained from $\Gamma^{\prime}\Rightarrow\alpha$ by ($\neg\Rightarrow$). By (IH), $\Gamma^{\prime}\models_{k}\alpha$. Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma^{\prime}\cup\{\neg\alpha\}$. Then, $v\in\mathcal{L}_{k}^{IPL}$ such that $v\models\Gamma^{\prime}$ and so $v\models\alpha$, by hypothesis, which contradicts the fact that $v\models\neg\alpha$. From this, $v\not\models\Gamma^{\prime}\cup\{\neg\alpha\}$ for every $v\in\mathcal{L}_{k+1}^{IPL}$, which means that $\neg\alpha,\Gamma^{\prime}\models_{k+1}$.
($\Rightarrow\neg$): Assume that $\Gamma\Rightarrow\Delta=\Gamma\Rightarrow\neg\alpha$ is obtained from $\alpha,\Gamma\Rightarrow$ by ($\Rightarrow\neg$). By (IH), $\alpha,\Gamma\models_{k}$. Let $v\in\mathcal{L}_{k+1}^{IPL}$ such that $v\models\Gamma$. Suppose that $v(\neg\alpha)=\mathbf{F}$. By definition of $\mathcal{M}_{IPL}$, $v(\alpha)=\mathbf{T}$, hence $v\models\Gamma\cup\{\alpha\}$, which contradicts the fact that $\alpha,\Gamma\models_{k}$, since $v\in\mathcal{L}_{k}$. Now, suppose that $v(\neg\alpha)=\mathbf{U}$. By Definition 8.6 of level valuations, there exists $w\in\mathcal{L}_{k}^{IPL}$ such that $w(\neg\alpha)=\mathbf{F}$ and $w\models\Gamma$. As proven above, $w(\alpha)=\mathbf{T}$ and so $w\models\Gamma\cup\{\alpha\}$, contradicting the fact that $\alpha,\Gamma\models_{k}$. From this, it follows that $v(\neg\alpha)=\mathbf{T}$. That is, $v\models\Gamma\Rightarrow\neg\alpha$ for every $v\in\mathcal{L}_{k+1}^{IPL}$.
This concludes the proof. ∎
Theorem 8.12 (Soundness of LI w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
For every set $\Gamma\cup\{\varphi\}$ of formulas over $\Omega$: $\Gamma\vdash_{IPL}\varphi$ implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
Proof.
Suppose that $\Gamma\vdash_{IPL}\varphi$. By definition, there exists a finite set $\Gamma_{0}\subseteq\Gamma$ such that the sequent $\Gamma_{0}\Rightarrow\varphi$ is derivable in LI. But then, there exists $k\geq 0$ such that $\Gamma_{0}\vdash_{IPL}^{k}\varphi$. By Proposition 8.11, $\Gamma_{0}\models_{k}\varphi$. This implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$, given that $\mathcal{L}_{IPL}\subseteq\mathcal{L}_{k}^{IPL}$. ∎
We will concentrate now on the completeness proof of LI w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$.
@@
Recall from Remark 6.3 that, in any Tarskian and finitary logic L, if $\Gamma\nvdash_{\bf L}\varphi$ then there exists a $\varphi$-saturated set $\Delta$ in L such that $\Gamma\subseteq\Delta$. This property holds, in particular, in the logic generated by LI (that is, IPL). Moreover, it is easy to prove that, if $\Delta$ is a $\varphi$-saturated set in IPL then: $\alpha\in\Delta$ implies that $\neg\alpha\notin\Delta$; $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; $\alpha\to\beta\in\Delta$ and $\alpha\in\Delta$ implies that $\beta\in\Delta$; and $\beta\in\Delta$ implies that $\alpha\to\beta\in\Delta$.
@@
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by LI (that is, IPL). The valuation associated to $\Delta$ in IPL is the function $v_{\Delta}:For(\Omega)\to V_{IPL}$ defined inductively as follows (here, $p$ is a propositional variable and $\alpha,\beta\in For(\Omega)$):